Two waves are traveling in the same direction along a stretched string. The waves are 90.0
Answer #1

Writing down the equation for each wave
Answer #2

y1 = 4.0 sin (wt+90)
y2 =.0 sin (wt)
these waves are 90 deg out of phase
at superposition > resultant amplitude (R)
R = y1+y2 = 4.0 sin (wt+90) + 4.0 sin (wt)
R = 4.0 [sin (wt+90) + sin (wt) ]
--------------------
sin c + sin d = 2 sin (c+d)/2 cos (c-d)/2
sin (wt+90) + sin (wt) = 2 sin [wt + 45] cos [45]
------------------------------
R = {4.0 * 2 cos (45)} sin (wt+45)
R = A sin (wt+ phi)
A = amplitude = 4.0 * [2 cos (45)]
A = 4.0 * [2 cos (45)]
A = 4 *1.414 = 5.656